OpenStudy (anonymous):

The maker of an automobile advertises that it takes 12 seconds to accelerate from 20 kilometers per hour to 65 kilometers per hour. Assuming constant acceleration, compute the acceleration in meters per second per second. Round your answer to three decimal places.

OpenStudy (anonymous):

It's just average rate of change.

OpenStudy (anonymous):

\[ \frac{\Delta v}{\Delta t} \]

OpenStudy (anonymous):

A. 1.968 m/sec2 B. 1.042 m/sec2 C. 1.250 m/sec2 D. 2.344 m/sec2 E. 2.361 m/sec2

OpenStudy (anonymous):

those are the options...

OpenStudy (anonymous):

helllooooo

OpenStudy (anonymous):

(delta v)/(delta t)=(65-20)km/hr*(1000 m/km)*(1 hr/3600 s)*(1/12 s)=1.042 m/s

OpenStudy (anonymous):

1) you find the change in velocity 2) convert it to m/s 3) divide by the change in time to get acceleration

OpenStudy (anonymous):

thank you!!